OpenStudy (anonymous):

okay..still don't understand these...solve the nonlinear system of equations {x^2+y^2=16 {x+y=4

OpenStudy (hba):

ok u take the 2nd eqyation that is x+y=4 take the value of x and plug it in the first one

hero (hero):

That's one way to do it

OpenStudy (anonymous):

is there another way?...one that's really easy perhaps? lol

hero (hero):

There's another way, but it isn't necessarily "easier"

OpenStudy (shane_b):

There are several ways: gaussian elimination, substitution, matrices....but I think substitution (as shown above) is pretty easy with simple ones like this.

OpenStudy (anonymous):

would anyone mind writing out some of the work?....not the answer but the beginning of the work showing the substitution

OpenStudy (shane_b):

Look at the second equation and solve for say, y: y=4-x Plug that into the first equation and solve for x. Once you know the value of x, you can use one of the original equations to solve for y,\.

OpenStudy (shane_b):

Plugging it in as I mentioned would look like this:\[x^2+(4-x)^2=16\]

hero (hero):

{x^2+y^2=16 {x+y=4 1. Square both sides on the second equation: \((x+y)^2 = 4^2 \) 2. Expand the left and right sides: \(x^2 + 2xy + y^2 = 16\) 3. Set 16 = 16 \(x^2 + y^2 = x^2 + 2xy + y^2\) subtract \(x^2\) and \(y^2\) from both sides: \(0 = 2xy\) 5. Divide both sides by 2: \(0 = xy\) 6. Therefore, (x,y) = (4,0) (x,y) = (0,4)

OpenStudy (anonymous):

oooo okay i get it now....thanks so much for all of your help!!!

hero (hero):

That's just an approach. Not saying it's the most logical one.

OpenStudy (anonymous):

well it makes sense...it was what my teacher was trying (and failing) to teach us hahaha

hero (hero):

I don't think your teacher taught you this way. Probably some other method.

hero (hero):

Do you have any more of these?

OpenStudy (anonymous):

ya i have 2 more

hero (hero):

Let me see them

OpenStudy (anonymous):

well this one is slightly different... 1. solve the system to determine the number of real solutions {x^2+y^2=16 {4x^2-9y^2=36 2. solve.... {x^2+y^2=63 {x^2-3y^2=27

OpenStudy (anonymous):

okay...i think i got it, thanks again for your help Hero!

hero (hero):

Actually, it would probably be better if you solved for y^2 and set y^2 = y^2

hero (hero):

For example for the first one, you should get \(y^2 = y^2\) \(16-x^2 = \frac{4x^2 - 36}{9}\)

hero (hero):

Then you can just cross multiply to get \(9(16 - x^2) = 4x^2 - 36\)

hero (hero):

\(144 - 9x^2 = 4x^2 - 36\) \(144 +36 = 4x^2 + 9x^2\) \(180 = 13x^2\) \(\large\frac{180}{13} = x^2\)

hero (hero):

You'd have to square root both sides to have x

hero (hero):

then substitute x back in to find y

hero (hero):

That's by far the simplest way to do it

hero (hero):

I hope that makes sense

OpenStudy (anonymous):

lol took me a minute to figure out what your were doing but ya now i see..it makes sense...a lot more sense than my teacher was making haha

hero (hero):

Well, at least you were able to understand it

hero (hero):

Do you want me to show you my exact steps in vyew?